The hypotenuse of a right triangle is 13m long. The longer leg is 7m longer than the shorter leg. Find the side lengths of the triangle.
Answer by richwmiller(17219)

5,12,13 is a famous right triangle
12-7=5
So it is this famous right triangle
13^2=x^2+(x+7)^2
x=5
x+7=12